OpenStudy (anonymous):

The students at Midtown Middle school sold flowers as a fundraiser in September and October. In October, they charged $1.50 for each flower. The October price was a 20% increase of the September price. Part A What was the price of the flowers in September? Enter your answer in the box. $ Part B The seventh-grade class earned 40% of the selling price of each flower. In September, they sold 900 flowers. In October, they sold 700 flowers. Select a choice from each drop-down menu to make a true statement. The seventh-grade class earned more money in

OpenStudy (anonymous):

help ill give fan and medal

OpenStudy (anonymous):

@IMStuck can u help

OpenStudy (texaschic101):

Oct -- 1.50 per flower Sept -- 20% of 1.50 = 0.2(1.5) = 0.30 : 1.50 - 1.30 = 1.20 per flower Part B : Sept : sold 900 flowers at 1.20 .....900 x 1.20 = 1080 40% of 1080 = (0.4)(1080) = 432 Oct : sold 700 flowers at 1.50...700 x 1.50 = 1050 40% of 1050 = (0.4)(1050) = 420 earned more money in Sept.

OpenStudy (imstuck):

Sorry, I got called away to tutor on tutor.com. Can't say no to that...it's my paying job! So sorry, but glad to see @texaschic101 helped!

OpenStudy (anonymous):

actually that is wrong the first part of the question is 1.25.